OpenStudy (anonymous):

cost function :the weekly marginal cost of producing x pairs of tenis shoes is c'(x)=12+500/x+1 where c(x) is cost in dollars if the fixed costs are 2000 per week find the cost function what is average cost pe pair of shoes if 1000pairs of shoes are produced each week

OpenStudy (dumbcow):

c(x) = 12x +500ln(x+1) + 2000 avg cost = c(1000)/1000 = 17.45